OpenStudy (anonymous):

3 3√3 one of the 3 is above the square root sign

OpenStudy (anonymous):

\[3\sqrt[3]{3}\]

OpenStudy (anonymous):

\[9 \sqrt{3} \approx 15.59\]

OpenStudy (anonymous):

no one of the threes is above the square root sign

OpenStudy (anonymous):

My bad.. \[3\sqrt[3]{3} = \sqrt[3]{3*3^{3}} = \sqrt[3]{81} \approx 4.326\]

OpenStudy (anonymous):

thanks